(1/4-1/x^2)/(1/2+1/x) can someone help me simplify this complex fraction?
Answer by mukhopadhyay(490)

(1/4-1/x^2)/(1/2+1/x)
(1/4-1/x^2) = [(1/2)^2 - (1/x)^2]
= (1/2 + 1/x)(1/2 - 1/x) ... (1)
When (1) is divided by (1/2 + 1/x):
(1/2 + 1/x)(1/2 - 1/x)/(1/2 + 1/x) = (1/2 - 1/x)